OpenStudy (anonymous):

44=5x-11 solve for x

OpenStudy (anonymous):

5 x 11 = 55 55 - 11 = 44

OpenStudy (anonymous):

@satellite73 is that okay

OpenStudy (anonymous):

\[44=5x-11\] \[44+11=5x-11+11\] \[55=5x\] \[55\div5=5x\div5\] \[11=x\]

OpenStudy (anonymous):

44+11=5x 55=5x x=11

OpenStudy (anonymous):

@Rohangrr sure you solved it in your head and proved that you were right

OpenStudy (anonymous):

lolxxx